If a third degree polynomial has a line x-intercept at x=a, discuss what this implies about the linear and quadratic factors of that polynomial


1
Expert's answer
2021-10-14T18:04:01-0400

Taking into account that a third degree polynomial "f" has alone "x"-intercept at "x=a", it has a unique root "x=a." If the multiplicity of this root is 3, then "f(x)=b(x-a)^3," and in this case the polynomial has three linear factors. If the root "x=a" has multiplicity 2, then it has linear factor "x-c" different from "x-a" , and polinomial has the root "x=c\ne a," which is impossible according to uniqueness of a root. Thererfore, "x=a" can not be of multiplicity 2. If the multiplicity of this root is 1, then "f(x)=(x-a)(mx^2+nx+k)," and the polynomial "h(x)=mx^2+nx+k" has no roots. It follows that the polynomial "g" is irreducible, and hence it is the quadratic factor of the polynomial "f".